OpenStudy (anonymous):

I am looking for solution step by step for this integral, I know what the answer is. \[\int\limits_{0}^{\infty}\frac{x ^{3}}{e ^{x}-1}dx=\frac{\pi ^{4}}{15}\]

OpenStudy (jamesj):

*bookmark. Going to bed now, but I'll write this out for you tomorrow. Nice question.

OpenStudy (anonymous):

tell me about it, it is really hard..

OpenStudy (jamesj):

Ok... \[ \int_0^\infty \frac{x^3}{e^x - 1} dx = \int_0^\infty \frac{x^3e^{-x}}{1-e^{-x}} \] \[ = \int_0^\infty \sum_{n=0}^\infty x^3 e^{-(n+1)x} dx \] \[ = \sum_{n=0}^\infty \int_0^\infty x^3 e^{-(n+1)x} dx \] Changing variable \( u = (n+1)x \), this is equal to \[ \sum_{n=0}^\infty \int_0^\infty \frac{u^3}{(n+1)^4} e^{-u} du \] \[ = \sum_{n=0}^\infty \frac{1}{(n+1)^4} \Gamma(4) \ \hbox{, by definition of the Gamma function }\] \[ = 6 \ \sum_{n=1}^\infty \frac{1}{n^4} \ \ \ \ \ \hbox{, as } \Gamma(4) = 3! = 6 \] \[ = 6 \ \frac{\pi^4}{90} \ = \ \frac{\pi^4}{15} \]

OpenStudy (jamesj):

We can now generalize this to \[ \int_0^\infty \frac{x^t}{e^x - 1} dx = \Gamma(t+1)\zeta(t+1) \]

OpenStudy (anonymous):

actually, my question contains \[\zeta(s)=\sum_{n=1}^{\infty}\frac{1}{n ^{s}}\] \[\zeta(4)=1+\frac1{2^{4}}+\frac1{3^{4}}+\frac1{3^{4}}...=\frac {\pi ^{4}}{90}\] how it is solution..

OpenStudy (anonymous):

I did not understand how passed this step \[= \int\limits_0^\infty \sum_{n=0}^\infty x^3 e^{-(n+1)x} dx\]

OpenStudy (jamesj):

\[ \frac{1}{1-x} = 1 + x + x^2 + x^3 + ... \] provided |x| < 1. As for \( x \in (0, \infty) \) we have \( 0 < e^{-x} < 1 \), \[ \frac{e^{-x}}{1-e^{-x}} = e^{-x} ( 1 + e^{-x} + e^{-2x} + e^{-3x} + ...) \] \[ = e^{-x} + e^{-2x} + e^{-3x} + e^{-4x} + ... \]

OpenStudy (jamesj):

does this make sense now?

OpenStudy (anonymous):

yes it does, I am still waiting for \[\zeta(4)=\frac {\pi^4}{90}\]

OpenStudy (jamesj):

I see. Let me try and find one of the more elementary proofs.